# Truck Parking Usage Prediction with Decomposed Graph Neural Networks

### v2

@@ Abstract
Truck parking on freight corridors faces the major challenge of insufficient parking spaces. This is exacerbated by the Hour-of-Service (HOS) regulations, which often result in unauthorized parking practices, causing safety concerns. It has been shown that providing accurate parking usage prediction can be a cost-effective solution to reduce unsafe parking practices. In light of this, existing studies have developed various methods to predict the usage of a truck parking site and have demonstrated satisfactory accuracy. However, these studies focus on a single parking site, and few approaches have been proposed to predict the usage of multiple truck parking sites considering spatio-temporal dependencies, due to the lack of data. This paper aims to fill this gap and presents the Regional Temporal Graph Neural Network (RegT-GCN) to predict parking usage across the entire state to provide more comprehensive truck parking information. The framework leverages the topological structures of truck parking site locations and historical parking data to predict the occupancy rate considering spatio-temporal dependencies across a state. To achieve this, we introduce a Regional Decomposition approach, which effectively captures the geographical characteristics of the truck parking locations and their spatial correlations. Evaluation results demonstrate that the proposed model outperforms other baseline models, improving performance by more than .
@@
The literature identifies increasing truck drivers’ accessibility to real-time parking usage information [3] and drivers’ perceptual interpretations [4, 5] as major factors influencing such parking challenges. According to the survey studies in the region [6], members of the Mid-America Association of State Transportation Officials (MAASTO) from eight states, Indiana, Iowa, Kansas, Kentucky, Michigan, Minnesota, Ohio, and Wisconsin, have collaborated to build a real-time multistate Truck Parking Information Management System (TPIMS) to provide more practical truck parking information to truck drivers [7, 8]. TPIMS, fully operational since January 4, 2019, provides real-time parking usage information to truck drivers through dynamic message signs, smartphone applications, traveler information websites (e.g. 511 traveler information) and other forms.
@@
Since 2019, the research team has developed the TPIMS data archiving system (TPIMS-DAS), designed to store effectively and track truck parking information [11]. Leveraging the substantial benefits derived from an abundance of comprehensive and detailed information on truck parking sites, users can extract and aggregate truck parking data from eight states in MAASTO using TPIMS-DAS. It also provides a sufficient amount of truck parking data to analyze truck parking usage patterns [12] and train reliable models. In this research, we aim to develop a data-driven, predictive model for future parking usage, grounded in spatio-temporal dependencies manifested by geographical locations of parking sites on highway corridors and their historical usage patterns.
To this end, we propose a Regional Temporal Graph Neural Network (RegT-GCN), comprising a Graph Neural Network with temporal capability and a novel decomposition module, Regional Decomposition, to capture regional/spatial relationships (Figure 1). In this study, a region is defined as a state to incorporate their unique characteristics (e.g., state laws). The significant historical parking data allows the model to capture temporal dependencies of multi-state-wide truck parking usage to generate more accurate prediction results. We conduct several experiments to evaluate the model against various state-of-the-art models and for the importance of Regional Decomposition for Graph Neural Networks.
@@
A spatio-temporal model to predict the usage of multiple truck parking sites.
A novel Regional Decomposition method, which leverages the regional relationships to create subgraphs for each region (i.e., state) and effectively captures the spatio-temporal dependencies.
The comprehensive truck parking data, which is aggregated across multi-states to train their spatio-temporal dependencies and evaluate the prediction accuracy of occupancy rates.
Extensive quantitative evaluations, which demonstrate that our RegT-GCN significantly outperforms state-of-the-art baseline models.
@@ II Literature Review
Parking prediction methods can be classified into two main categories: urban parking prediction, which is related to parking prediction within cities, and truck parking prediction, which deals with predicting parking usage for trucks. Typically, urban parking prediction methods offer a shorter prediction horizon, ranging from 5 to 15 minutes ahead. In contrast, truck parking prediction provides predictions for up to 1 hour ahead [10]. Regardless of the distinctions in time frames between these parking prediction tasks, it should be noted that both tasks often share fundamental concepts and are devised with a broader perspective that encompasses diverse parking scenarios.
II-A Statistical Prediction Model for Truck Parking Usage
Bayraktar et al. [13] developed an intelligent parking-management system for commercial vehicles in Florida, analyzed data, and created a prediction model for rest area occupancy with a Kalman filter. As the Kalman filter prediction-correction mechanism can be utilized with historical data and adapted to the online status, its model achieved better performance. Sadek et al. [9] identified several patterns such as the regular, surplus, discharge, holiday, and weekend patterns. They proposed two models to predict future usage with lower RMSE. One is the hybrid Fourier model of the Trend Switching model, which predicts usage patterns starting from the regular pattern and switches to surplus or discharge later. Another model, named the Trend Shifting model, shifts the existing pattern to match the observed data. Haque et al. [14] compared four different Poisson-related models and stated that the Poisson model with propensity and threshold-specific constant is the best. The following non-homogeneous Poisson model proposed in [15] surpasses the other methods, such as the multivariate spatio-temporal model [16] curve similarity and historical average. Moreover, the queuing model has a non-homogeneous arrival rate and time-varying service time distribution. Hamidreza et al. [17] presented a queuing model in which arrivals follow a non-homogeneous Poisson process and predicted parking usage probabilistically from microscopic perspectives of each parking and macroscopic perspectives of aggregated parking information.
II-B Machine Learning Prediction Model for General Parking Usage
Although these statistical models successfully predicted with reasonable accuracy, the models depend on sequential data such as parking availability and arrival and departure rates. Recent researchers have used machine learning techniques to integrate all related features such as weather conditions, daily patterns, and parking location and predict future usages with better precision [18, 19, 10]. For the research on citywide parking, Zhao and Zhang [18] developed six types of predicting models (FM1, FM2, linear regression, support vector machine (SVM), backpropagation neural network, and autoregressive integrated moving average). SVM and its ensemble model achieved lower RMSE as it is easier to find the global optimum. Zheng et al. [19] first predicted the distribution of the parking arrival and departure patterns with the Markov birth-and-death process. Then, with parameter estimation, they predicted the parking demand for the short term. Hao et al. [10] analyzed the parking usage pattern with data aggregation at several parking sites. They found that there are tendencies in the on-working period (Sunday night to Friday daytime) and off-working period (others), in which the usage remains lower. They also developed spatio-temporal models with attention modules on the decoder and achieved high-precision results. Lastly, Low et al. [20] interestingly focused on predicting the duration of commercial vehicle parking. Before putting the data into the regression models, they generated the incomplete field with generative adversarial multiple imputation networks (GAMIN), then output the predicted dwell time.
II-C Spatio-Temporal Prediction Model for General Parking Usage
@@
Most prior research on predicted truck parking utilization focuses primarily on historical parking usage data and static variables, such as speed limit and annual average daily traffic [10]. However, it overlooks the inclusion of other significant predictors. According to Boris and Brewster [1], the key determinants guiding truck drivers’ decision-making process are parking location and the quality of the available amenities. Consequently, integrating topological structures into the predictive model is justified, given its potential to account for spatial variations in the influential factors.
@@
Graph Neural Networks (GNN) [21] are well known for understanding and representing the topological structures of the input. GNN has been widely used in the transportation domain for its ability to capture spatial dependencies. A road network is naturally functional as a non-Euclidean graph structure [22]. Yang et al. [23] first leveraged Graph-Convolutional Neural Networks (GCN) on urban parking to extract the spatial dependencies of traffic flow in large-scale networks and achieved a low Mean Absolute Error (MAE) in Pittsburgh downtown. They also incorporated Recurrent Neural Networks (RNN) and Long Short Term Memory (LSTM) to capture temporal dependencies with additional related features such as weather data and traffic speed to achieve the parking usage prediction with regard to spatio-temporal dependencies. Zhang et al. [24] proposed a hierarchical graph convolution structure. They addressed the scarcity of real-time parking availability information with a semi-supervised approach and spatio-temporal dependencies among parking lots. Despite their success in urban parking with spatio-temporal dependencies, they have yet to conduct modeling and experiments on truck parking usage prediction due to the lack of sufficient data.
Compared to approaches applied to urban car parking, truck parking is more relative to truck drivers’ route decisions, and critical characteristics such as the existence of amenities, capacity limitations, and diurnal patterns directly affect the predicted results [9]. The parking selection strategies of drivers exhibit distinction, with truck drivers typically opting for parking site along highways while car drivers predominantly seek parking proximate to their urban destinations. Furthermore, most truck drivers routinely work and find their resting place during night or weekends, different from the temporal patterns observed in urban parking scenarios, where drivers look for parking during the day [25].
This study aims to predict future parking usage patterns considering the highway network’s topological structures and spatio-temporal dependencies of the network. Specifically, we develop a graph convolution model to capture these complex characteristics, enabled by extensive data on highway networks and parking characteristics, such as the number of amenities and the distance from the nearest city.
@@ III-A Problem Formulation
In order to capture the spatio-temporal dependencies, this research first engages in understanding regional relationships across truck parking sites. Based on our preliminary analysis [12] and insights from previous studies on the efficacy of edge reduction techniques [19], [26], we assume that the regional relationship within the truck parking network can be leveraged to reduce graph complexity. We hypothesize that a regionally decomposed graph structure, as opposed to one connected graph, offers advantages in terms of both predictive performance and computational efficiency.”
In this paper, we define small partitions in the truck parking sites as an independent region and explore the spatial dependencies among truck parking sites: state. Site-specific parking behavior deliberates region selection to be grounded in practicality and the unique characteristics exhibited by states [26]. For example, freight plans vary by state, so they recommend different values and strategies for developing, funding, and maintaining shared facilities. It is also beneficial to consider the state as a region from a data aggregation perspective. Hence, our regional selection prioritizes interpretability and practical outcomes.
@@ III-B Regional Temporal Graph Convolutional Network
In this paper, we built the Graph Neural Network, which incorporates the geographical characteristics in nodes integrating spatial and temporal aspects with regional graph input. Our model comprises two main components: a Graph Convolution Network (GCN), inspired by the Attention Temporal Graph Convolutional Network (A3TGCN) [27], as a spatial module and a recurrent unit as a temporal module. A structual GCN handles the node embeddings of the truck parking location graph, and a spatial GCN compresses the spatial dependencies between sites taken the distances between each pair of truck parking sites. Given the sequence of input timestamps , the temporal model predicts site occupancy rates up to future time steps.
@@ IV-A Data Collection and Pre-processing
A collaborative effort among eight states in MAASTO led to the development of a real-time multi-state Truck Parking Information Management System (TPIMS). Operational since 2019, the TPIMS data archival system (TPIMS-DAS) has archived more than 100 million records of truck parking information (up to October 2023).
Figure 3 shows all truck parking sites in MAASTO states and where we have collected parking data. Currently, TPIMS-DAS has archived data from 144 parking sites (public and private) of the eight participating MAASTO states. The number of current sites reporting data in each state is Iowa: 44, Illinois: 19, Kansas: 18, Kentucky: 13, Michigan: 14, Minnesota: 7, Ohio: 18, and Wisconsin: 11. About 160,000 parking records are archived in TPIMS-DAS daily; hence more than 1000 records per site on average. Due to maintenance periods in some sites and a lack of consequent data, we picked 105 sites for the graph construction and usage prediction.
Before applying our model, we handled the raw data in the appropriate format. We extracted dependent features from archived data following the formats in Table I. Week ID, day ID, and the hour ID are dynamically changed, and other features are statically decided. These features were identity encoded and used as the input features of the nodes. The occupancy rate is calculated using the available data and capacity, and is the target feature of the model.
Edges are composed of two different sites, and with the characteristics of road networks, not all nodes are connected. To realize that, we selected sites within 40 miles, which is approximately a 35-minute driving distance, and connected them for a single connected graph and regional subgraphs. Each edge has the actual driving distance as weights extracted from Bing Maps Locations API. For this research, we used 10-minute frequency data, while some sites have different data collection frequencies. Therefore, we interpolated the occupancy rate data, taking the average of two consecutive values to fill in the blank data. Other dynamic features were determined to be the same number as previous values.
@@
| Amenity | The number of amenities (e.g., restrooms, handicap accessible, Wifi) |
@@ IV-C Model Comparison
We compared our proposed framework, Regional Temporal GCN (RegT-GCN), with the following baseline methods: Stacked GRU, Stacked GCN [29], Temporal GCN (T-GCN) [27], Temporal Graph Convolution LSTM (LTGC) [30] and Graph SAGE [31]. They were all applied RMSProp and the learning rate is with the decaying rate of . Input channels are 8 for all models and output channels are 512 for LSTM, T-GCN, and LTGC and 256 for the other models. For this experiment, we used one GeForce RTX 3080 for training and inference. Experimentally, our model takes around 15 minutes for training and 1 minute for inference.
@@
T-GCN is composed of GRU as a temporal module with attention-based aggregation named Attention Temporal GCN (A3T-GCN) on its spatial module [27].
@@
Graph SAGE is a general framework for inductive representation learning on large graphs with low-dimensional node embeddings.
RegT-GCN uses A3T-GCN for the spatial module and GRU for the temporal module. Input subgraphs are created with Regional Decomposition.
@@ IV-D Truck Parking Usage Prediction
Each model was trained on of the data (March 1st to March 3rd, 2022) and tested on the rest of of the data (March 3rd to March 14th, 20202). Table II demonstrates the performance comparison with baseline models. Comparing the temporal and spatial models, Stacked GCN predicted generally better than Stacked GRU, which indicates the spatial model learns better representations of truck parking sites than the temporal model. In comparison, most spatio-temporal models improved the model performance, especially in shorter time horizons compared to Stacked GRU and Stacked GCN. Across the models, our RegT-GCN outperforms the baselines. Specifically in longitudinal analysis, Stacked GRU decreased its performance compared to Stacked GCN since the temporal model highly depends on the predicted time horizons. At longer time horizons, the result of Stacked GCN predicted a better or closer accuracy to spatio-temporal models, while Stacked GRU and T-GCN performance dropped the prediction accuracy. Across most time horizons, our RegT-GCN successfully demonstrates improvements in truck parking prediction performance.

### v3

@@ Abstract
Truck parking on freight corridors faces the major challenge of insufficient parking spaces. This is exacerbated by the Hour-of-Service (HOS) regulations, which often result in unauthorized parking practices, causing safety concerns. It has been shown that providing accurate parking usage prediction can be a cost-effective solution to reduce unsafe parking practices. In light of this, existing studies have developed various methods to predict the usage of a truck parking site and have demonstrated satisfactory accuracy. However, these studies focused on a single parking site, and few approaches have been proposed to predict the usage of multiple truck parking sites considering spatio-temporal dependencies, due to the lack of data. This paper aims to fill this gap and presents the Regional Temporal Graph Convolutional Network (RegT-GCN) to predict parking usage across the entire state to provide more comprehensive truck parking information. The framework leverages the topological structures of truck parking site locations and historical parking data to predict the occupancy rate considering spatio-temporal dependencies across a state. To achieve this, we introduce a Regional Decomposition approach, which effectively captures the geographical characteristics of the truck parking locations and their spatial correlations. Evaluation results demonstrate that the proposed model outperforms other baseline models, showing the effectiveness of our regional decomposition. The code is available at https://github.com/raynbowy23/RegT-GCN.
@@
The literature identifies increasing truck drivers’ accessibility to real-time parking usage information [3] and drivers’ perceptual interpretations [4, 5] as major factors influencing such parking challenges. Along with the survey studies in the region [6], members of the Mid-America Association of State Transportation Officials (MAASTO) from eight states, Indiana, Iowa, Kansas, Kentucky, Michigan, Minnesota, Ohio, and Wisconsin, have collaborated to build a real-time multistate Truck Parking Information Management System (TPIMS) to provide more practical truck parking information to truck drivers [7, 8]. TPIMS, fully operational since January 4, 2019, provides real-time parking usage information to truck drivers through dynamic message signs, smartphone applications, traveler information websites (e.g. 511 traveler information) and other forms.
@@
To this end, we aim to develop a data-driven, predictive model for future parking usage, grounded in spatio-temporal dependencies of parking sites on highway corridors and their historical usage patterns. In this research, we propose a Regional Temporal Graph Convolutional Network (RegT-GCN), comprising a Graph Convolutional Network (GCN) with temporal capability and a novel decomposition module, Regional Decomposition, to capture regional and spatial relationships (Figure 1). In this study, a region is defined as a state to incorporate their unique characteristics (e.g., state laws). The significant historical parking data allows the model to capture temporal dependencies of multi-state-wide truck parking usage to generate more accurate prediction results. We conduct several experiments to evaluate the model against various models and raised the importance of Regional Decomposition technique for Graph Convolutional Networks.
@@
We propose a spatio-temporal model to predict the usage of multiple truck parking sites.
Our novel Regional Decomposition method leverages the regional relationships to create subgraphs for each region (i.e., state) and effectively captures the spatio-temporal dependencies.
We construct the comprehensive truck parking dataset, which is aggregated across multi-states to train their spatio-temporal dependencies and evaluate the prediction accuracy of occupancy rates.
We conduct extensive quantitative evaluations including temporal predictions, spatial analysis, and sensitivity analysis, which demonstrate that our RegT-GCN significantly outperforms baseline models.
@@ II Literature Review
II-A Spatio-Temporal Prediction in Transportation
Spatio-temporal dependencies in the transportation field are notably effective for traffic flow prediction as they take advantage of highway network topology. Those approaches adopted graph convolutions [11, 12, 13], attention mechanism [14, 15], and transformer [16]. Specifically, GCNs [17] are highly effective in capturing spatial embeddings used with temporal embedding modules [11, 13] while the attention mechanism improves its performance in long-range temporal dependencies working in conjunction with spatial embedding models [14, 16]. Recent works such as STID [18] and STAEformer [19] extended spatial-temporal dependencies with more sophisticated embedding networks. All these methods, however, lack a deep investigation of spatial embeddings that align with urban and truck parking usage prediction.
II-B Parking Usage Prediction
Parking prediction methods can be classified into two main categories: urban and truck parking prediction. Urban parking prediction typically offers a shorter horizon, ranging from 5 to 15 minutes, while truck parking prediction extends more than an hour ahead [10]. Despite these differences in time frames, both tasks share underlying concepts and are designed to address a variety of parking scenarios. Models that operate parking prediction combine historical data and real-time updates [20, 21]. This is extended in the pattern recognition approach [9, 22]. Vital et al. [23] further enhanced prediction accuracy by capturing spatio-temporal dependencies. Recent studies [11, 24, 25] have adopted machine learning techniques to integrate additional features, such as weather, daily patterns, and parking location. Although feature embeddings are effective for parking prediction [10], these studies limit the prediction capability within a few sites.
II-C Spatio-Temporal Representation in Parking Prediction
@@
While most prior research on predicting truck parking utilization focuses primarily on historical parking usage data and static variables, it overlooked the inclusion of spatial dependencies as the key determinants of truck driver’s decision-making process [1]. This gap highlights the need for integrating topological structures into predictive models to account for spatial variations in influential factors.
@@
This need for spatial integration aligns with the road network’s inherent structure, which naturally functions as a non-Euclidean graph [26]. Yang et al. [27] first leveraged GCN on urban parking to extract the spatial dependencies of traffic flow in large-scale networks and incorporated Recurrent Neural Networks (RNN) and Long Short Term Memory (LSTM) to capture temporal dependencies with additional related features. Zhang et al. [28] addressed the scarcity of real-time parking availability information with a semi-supervised approach and spatio-temporal dependencies among parking lots. Despite their success in urban parking with spatio-temporal dependencies, they have yet to conduct modeling and experiments on truck parking usage prediction due to the lack of sufficient data.
Therefore, our method benefits from the model with spatio-temporal embeddings while adapting truck parking unique features such as amenities, capacity limits [9], and diurnal patterns [29]. In contrast to the introduction of the spatio-temporal prediction model in an urban parking scenario [30], we consider the unique regional dependencies to capture a large truck parking network.
@@ III-A Problem Formulation
In order to capture the spatio-temporal dependencies, this research first engages in understanding regional relationships across truck parking sites. Based on our preliminary analysis [31] and insights from previous studies on the efficacy of edge reduction techniques [32, 33], we assume that the regional relationship within the truck parking network can be leveraged to reduce graph complexity. We hypothesize that a regionally decomposed graph structure, as opposed to one connected graph, offers advantages in terms of both predictive performance and computational efficiency.
In this paper, we define small partitions in the truck parking sites as an independent region and explore the spatial dependencies among truck parking sites. Site-specific parking behavior deliberates region selection to be grounded in practicality and the unique characteristics exhibited by states [34]. For example, freight plans vary by state, so they recommend different values and strategies for developing, funding, and maintaining shared facilities. It is also beneficial to consider the state as a region from a data aggregation perspective. Hence, our regional selection prioritizes interpretability and practical outcomes.
@@ III-B Regional Temporal Graph Convolutional Network
In this paper, we built the graph-based spatio-temporal model, which incorporates the geographical characteristics in nodes integrating spatial and temporal aspects with regional graph input. Our model comprises two main components: a GCN, inspired by the Attention Temporal Graph Convolutional Network (A3TGCN) [35], as a spatial module and a recurrent unit as a temporal module. A structual GCN handles the node embeddings of the truck parking location graph, and a spatial GCN compresses the spatial dependencies between sites taken the distances between each pair of truck parking sites. Given the sequence of input timestamps , the temporal model predicts site occupancy rates up to future time steps.
@@ IV-A Data Collection and Pre-processing
Figure 3 (Left) shows all truck parking sites in MAASTO states and where we have collected parking data. Currently (as of Oct, 2023), TPIMS-DAS has archived data from 144 parking sites (public and private) of the eight participating MAASTO states. The number of current sites reporting data in each state is Iowa: 44, Illinois: 19, Kansas: 18, Kentucky: 13, Michigan: 14, Minnesota: 7, Ohio: 18, and Wisconsin: 11. About 160,000 parking records are archived in TPIMS-DAS daily; hence more than 1000 records per site on average. Due to maintenance periods in some sites and a lack of consequent data, we picked 104 sites for the graph construction and usage prediction. Right figures on Figure 3 show average occupancy rate in daily and monthly from our previous research [31].
To apply TPIMS data in our model, we pre-processed the raw data in the appropriate format. We extracted dependent features from archived data following the formats in Table I. Week ID, day ID, and hour ID are dynamically changed, and others are static features corresponding to each site. These features were identity encoded and used as the input features of the nodes. The occupancy rate is calculated using the available data and capacity and is the target feature of the model.
Edges connect two different sites, and nodes are partially connected with the characteristics of highway networks. Following that, we selected sites within 40 miles, which is approximately a 35-minute driving distance, then connected them to construct single connected graph and regional subgraphs. Each edge has the actual driving distance as weights extracted from Bing Maps Locations API. For this research, we used 10-minute frequency data, while some sites have different data collection frequencies. Therefore, we interpolated the occupancy rate data with the linear interpolation. Other dynamic features were determined to be the same number as the previous timestep. Please refer to more details from here (https://github.com/raynbowy23/TPIMSDataset).
@@
| Amenity | The number of amenities (e.g., restrooms, handicap accessible, Wi-Fi) |
@@ IV-C Model Comparison
We compared our proposed framework, Regional Temporal GCN (RegT-GCN), with the following baseline methods: Stacked GRU, Stacked GCN [37], Temporal GCN (T-GCN) [35], Temporal Graph Convolution LSTM (LTGC) [38], Graph SAGE [33], Spatial and Temporal Normalization (STNorm) [39], and Spatial Temporla Identity (STID) [18]. They were all applied RMSProp and the learning rate is with the decaying rate of . Input channels are 8 for all models and output channels are 512 for LSTM, T-GCN, and LTGC and 256 for the other models. For this experiment, we used one GeForce RTX 3080 for training and inference. Experimentally, our model takes around 15 minutes for training and 1 minute for inference.
@@
T-GCN is composed of GRU as a temporal module with attention-based aggregation (A3TGCN) on its spatial module [35].
@@
Graph SAGE [33] is a general framework for inductive representation learning on large graphs with low-dimensional node embeddings.
STID [18] integrates spatial and temporal identities to enhance predictive performance.
RegT-GCN uses A3TGCN for the spatial module and GRU for the temporal module. Input subgraphs are created with Regional Decomposition.
@@ IV-D Truck Parking Usage Prediction
Each model was trained on of the data (March 1st to March 3rd, 2022) and tested on the rest of of the data (March 3rd to March 14th, 2022). Table II demonstrates the performance comparison with baseline models. Comparing the temporal and spatial models, Stacked GCN predicted generally better than Stacked GRU, which indicates the spatial model learns better representations of truck parking sites than the temporal model. In comparison, most spatio-temporal models improved the model performance, especially in shorter time horizons compared to Stacked GRU and Stacked GCN. Across the models, our RegT-GCN outperforms the baselines. Specifically in longitudinal analysis, Stacked GRU decreased its performance compared to Stacked GCN since the temporal model highly depends on the predicted time horizons. At longer time horizons, the result of Stacked GCN predicted a better or closer accuracy to spatio-temporal models, while Stacked GRU, T-GCN, and STID performance dropped the prediction accuracy. Across most time horizons, our RegT-GCN successfully demonstrates improvements in truck parking prediction performance.